# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs

### v4

@@ Appendix A Contribution & Limitation
Despite the limited number of models, tasks, and judging types, our proposed systematic framework is agnostic to these criteria and can be used with any LLM judge. Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through position consistency and preference fairness. We also explore the factors influencing position bias and their quantitative impacts, such as the answer quality gap and lengths of prompt components. Particularly, our study is the first to establish a direct connection between quantified answer quality disparities and position bias, highlighting the urgent need for a comprehensive understanding of how position bias affects LLM judges. Additionally, the mutual agreement and disagreement analysis of LLM judges are insightful for identifying the difficulty of evaluation instances and beneficial for benchmark assessment and designs. All findings from our study enhance the understanding of the position bias of LLM-as-a-Judge, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
Social Benefits Our comprehensive understanding of position bias in LLM-as-a-Judge provides multiple benefits to the community. First, it enhances the reliability and trustworthiness of LLM judges in terms of repetition stability validation. Our focus and investigation on preference fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability. Moreover, our emphasized trade-offs between position consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work. Additionally, the baseline comparison analysis across judges and tasks offers reasonable judge model recommendations for more reliable, consistent, and fair evaluation results. Besides, the LLM agreement/disagreement analysis in Appendix.Sec.C 5 provides an applicable tool to identify the difficulty of judging certain instances. Together with findings from the significant impact of the answer quality gap on position bias, our study benefits future benchmark designs and evaluations and lays the groundwork for potential mitigation strategies of position bias. Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare Chen et al. (2024c), instruction following and prompting Zeng et al. (2023); Chua et al. (2024), multimodal assessment Chen et al. (2024a), and recommender systems Hou et al. (2024).
@@
Our framework of systematically understanding position bias has the following strengths: (1) scalability and ease of use - our proposed framework is scalable to various judges, models, tasks, and judging types. It is also straightforward and convenient to implement, facilitating a more comprehensive evaluation of LLM judges for future works. (2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, position consistency, and preference fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias. (3) accurate measurement - we formally define crucial metrics ($RS$ and $PF$) for evaluating LLM judges’ position bias that are underexplored by prior works and offer detailed explanations for their rationality and rigorousness.
@@
Appendix B Related Work
B.1 LLM-as-a-Judge
@@
In recent years, Large Language Models (LLMs) have emerged as a transformative technology, garnering global attention and stimulating substantial research into their applications. For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature (Zeng et al., 2023). However, it lacks scalability and reproducibility (Karpinska et al., 2021).
@@
As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks (Zheng et al., 2024b; Li et al., 2024; Chiang & Lee, 2023b; Liusie et al., 2024; Chen et al., 2023; Fu et al., 2023; Wang et al., 2023a; Kocmi & Federmann, 2023b; Ji et al., 2023; Liu et al., 2024; Zeng et al., 2023; Liu et al., 2023; Shen et al., 2023; Kocmi & Federmann, 2023a; Chiang & Lee, 2023a; Karpinska et al., 2021; Dubois et al., 2024; Kim et al., 2023; Liang et al., 2023; Chang et al., 2023; Zhong et al., 2023; Chia et al., 2023; Deb & Das, 2023; Ye et al., 2024; Min et al., 2023; Li et al., 2023d; c; b; Lu et al., 2024), including open-ended story generation (Chiang & Lee, 2023a), adversarial attacks (Chiang & Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi & Federmann, 2023a), and instruction following (Zeng et al., 2023).
@@
These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles (Zheng et al., 2024b; a; Pezeshkpour & Hruschka, 2023; Chiang & Lee, 2023b; Shen et al., 2023; Liusie et al., 2024; Qin et al., 2024; Chen et al., 2023; Huang et al., 2024; Zhu et al., 2023b; Raina et al., 2024; Chen et al., 2024b; Wang et al., 2023b; Liu et al., 2024; Zeng et al., 2023; Khan et al., 2024; Kim et al., 2023; Li et al., 2023b; d; Shi et al., 2024; Li et al., 2023c; Chen et al., 2024c; Chua et al., 2024; Chen et al., 2024a). As LLMs have made content generation significantly easier, the volume of generated responses has increased, making it impractical to rely solely on human evaluation. Therefore, cost-effective LLM judges are needed to assess these responses efficiently. LLM-as-a-Judge is typically employed in Q&A evaluation tasks, where the LLM judge is prompted to evaluate the quality of responses, usually generated by models, answering the questions.
@@
In many of these tasks, LLM judges have shown a high level of agreement with human evaluators (Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), yet in some tasks, they are less effective, largely due to inherent biases (Shen et al., 2023; Chiang & Lee, 2023b; Zheng et al., 2024a). Even in cases where agreement is high, judgments may still suffer from biases.
@@
When employing LLM-as-a-Judge, various types of judging are available, which can be categorized either by the scale or comparative method. From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment (Qin et al., 2024). By comparative method, it can be either score-based or relation-based (Li et al., 2023d).
@@
For example, pointwise scoring (Kim et al., 2023) lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics. Pairwise/Listwise Scoring (Karpinska et al., 2021; Zhu et al., 2023b; Chen et al., 2024a; Zheng et al., 2024b) prompts the LLM judges to score a pair/list of model-generated answers. Listwise Ranking is another relation-based assessment that, instead of giving a score, requires the LLM judge to rank a list of responses following some specified order (e.g., from best to worst). Pairwise comparative assessment (Karpinska et al., 2021; Chen et al., 2024a; Zheng et al., 2024b), on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
@@
B.2 Position Bias
Previous studies have discovered and investigated multiple types of biases, such as position bias (Zheng et al., 2024b; Li et al., 2024; Chen et al., 2024a; 2023; Chia et al., 2023; Chiang & Lee, 2023b; Khan et al., 2024; Zhu et al., 2023b; Zhong et al., 2023; Zheng et al., 2024a; Deb & Das, 2023; Zeng et al., 2023; Wang et al., 2023c; Qin et al., 2024; Wang et al., 2023b; Shi et al., 2024; Liusie et al., 2024; Liu et al., 2024; Kim et al., 2023; Karpinska et al., 2021; Hou et al., 2024; Kocmi & Federmann, 2023b; Pezeshkpour & Hruschka, 2023; Shen et al., 2023; Chiang & Lee, 2023a; Kocmi & Federmann, 2023a; Chen et al., 2024b; Li et al., 2023c; d; Raina et al., 2024), verbosity/length bias (Zheng et al., 2024b; Liu et al., 2024; Kim et al., 2023), self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b), selection bias (Zheng et al., 2024a; Pezeshkpour & Hruschka, 2023), and contextual bias (Liu et al., 2024; Zhou et al., 2024).
Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source (Li et al., 2023b; Chen et al., 2024c), proprietary commercial ones (OpenAI, 2023; Anthropic, 2024; Gemini Team, 2024), and fine-tuned models (Huang et al., 2024; Wang et al., 2023c; Zhu et al., 2023b; Li et al., 2023b; Kim et al., 2023).
To clarify, the aforementioned and following position bias refers to the concept within the context of LLM-as-a-Judge, meaning that LLM judge tends to favor responses based on their position in the prompt rather than their content. For example, in a pairwise comparative assessment scenario, if the LLM judge consistently selects the first response as superior even after switching the order of the two responses (same position, but different content), then a position bias occurs. Some other studies also use the term position bias (Ko et al., 2020; Aslanyan & Porwal, 2019; Blunch, 1984; Raghubir & Valenzuela, 2006), but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
Position bias is arguably the most prevalent and impactful bias among all. Chua et al. (2024) notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought (Wei et al., 2022) performance over self-training controls for all biases except position bias. Furthermore, Khan et al. Khan et al. (2024) point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
Moreover, there is ongoing debate over whether selection bias originates from position bias. Pezeshkpour & Hruschka (2023) argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity. Raina et al. (2024) contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
@@ B.2 Position Bias
B.3 Deal with Position Bias
Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts (Zheng et al., 2024b; Raina et al., 2024; Pezeshkpour & Hruschka, 2023; Zhao et al., 2021; Zhu et al., 2023a). Also, LLM judges are vulnerable to attacks (Raina et al., 2024; Chen et al., 2024b; Shi et al., 2024).
Parse data with position bias There are many ways to deal with the position bias. The naive way is to exclude the inconsistent judgments(i.e. if the LLM judge gives a positionally biased judgment on a pair or list of model responses) (Zheng et al., 2024b; Chen et al., 2024b; Wang et al., 2023b; Li et al., 2023b). While this ensures consistent and reliable remaining judgments, it does not resolve the fundamental problem. Moreover, if the LLM judge is highly biased, this method discards valuable evaluation instances and information, making it an ineffective and somewhat desperate measure.
To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging (Zheng et al., 2024b; Raina et al., 2024; Li et al., 2023b; Wang et al., 2023b) or regard inconsistency as a half-win or tie for relation-based judging (Zheng et al., 2024b; Li et al., 2023c) after swapping the order of model responses in the prompt. For instance, in a pairwise scoring scenario, if model A receives a score of 8 when put in the first position and 4 when put in the second position, the overall score for it compared to model B would be 6; in the pairwise comparative case, if model A wins when put in the first position but then lose when put in the second position, it counts as a tie or half-win for both model A and model B. The latter way of swapping + tie is proposed because intuitively the position bias is more likely to occur when the model responses get evaluated share a similar quality in terms of the evaluating metric. Our quantitative study also verifies this intuition, evidenced by the fact the LLM judge’s positional consistency (the percentage of positionally consistent pairs of judgment) is positively proportional to the answer quality gap between the model responses.
@@ B.3 Deal with Position Bias
Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping (Hou et al., 2024), split-and-merge (Li et al., 2023d), and multi-agent discussion (Li et al., 2023c; Khan et al., 2024). However, these methods are either costly and time-consuming (e.g., multi-agent discussion and review to reach agreement) or ineffective.
Furthermore,Liu et al. (2024) suggests that existing calibration techniques designed to reduce bias, including context decomposition (Li et al., 2023d), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a), and batch calibration (Zhou et al., 2024), are insufficient to align LLM evaluators, even with supervised data.
@@
For instance, Li et al. (2023d) proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%. However, the extraordinarily high consistency of the original evaluation may imply that the choice of answer pairs may be biased in terms of answer quality gap, a factor that our study proves to be significantly impactful. In other words, if the quality between the two model responses differs considerably, the LLM judges will make little position bias during judgment and hence is easy to calibrate. This example illustrates how lacking a comprehensive understanding of the factors affecting position bias can lead to overestimating or misjudging the effectiveness of methods proposed to resolve it and improve the performance of LLM-as-a-Judge.
@@
While prior research includes some positional preference results (Zheng et al., 2024b; Wang et al., 2023b; Zhu et al., 2023b; Li et al., 2023c; Liusie et al., 2024), none have conducted a detailed analysis. They focus primarily on positional consistency in relation to position bias, while we regard positional preference as an equally crucial component that needs to be examined, understood, and enhanced in future work.
@@
Apart from that, previous studies have yet to examine how stable LLM judges are across repetitions. MLLM-as-a-Judge (Chen et al., 2024a) conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments. In comparison, we investigate the repetition stability to determine whether position bias is solely influenced by the prompt’s positional information and the judge’s intrinsic properties or can be partly attributed to judgment randomness. If the judgments exhibit randomness over repetitions, then the position bias overlaps with repetition bias, complicating the issue and potentially invalidating previous findings.
B.4 Summary of Prior Work

### v5

@@ Appendix F Contribution & Limitation
Despite the limited number of models, tasks, and judging types, our proposed systematic framework is agnostic to these criteria and can be used with any LLM judge. Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through positional consistency and positional fairness. We also explore the factors influencing position bias and their quantitative impacts, such as the answer quality gap and lengths of prompt components. Particularly, our study is the first to establish a direct connection between quantified answer quality disparities and position bias, highlighting the urgent need for a comprehensive understanding of how position bias affects LLM judges. Additionally, the mutual agreement and disagreement analysis of LLM judges are insightful for identifying the difficulty of evaluation instances and beneficial for benchmark assessment and designs. All findings from our study enhance the understanding of the position bias of LLM-as-a-Judge, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
Social Benefits Our comprehensive understanding of position bias in LLM-as-a-Judge provides multiple benefits to the community. First, it enhances the reliability and trustworthiness of LLM judges in terms of repetition stability validation. Our focus and investigation on positional fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability. Moreover, our emphasized trade-offs between positional consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work. Additionally, the baseline comparison analysis across judges and tasks offers reasonable judge model recommendations for more reliable, consistent, and fair evaluation results. Besides, the LLM agreement/disagreement analysis in Appendix.5 provides an applicable tool to identify the difficulty of judging certain instances. Together with findings from the significant impact of the answer quality gap on position bias, our study benefits future benchmark designs and evaluations and lays the groundwork for potential mitigation strategies of position bias. Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare [18], instruction following and prompting [1, 9], multimodal assessment [19], and recommender systems [20].
@@
Our framework of systematically understanding position bias has the following strengths: (1) scalability and ease of use - our proposed framework is scalable to various judges, models, tasks, and judging types. It is also straightforward and convenient to implement, facilitating a more comprehensive evaluation of LLM judges for future works. (2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, positional consistency, and positional fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias. (3) accurate measurement - we formally define crucial metrics ($RS$ and $PF$) for evaluating LLM judges’ position bias that are underexplored by prior works and offer detailed explanations for their rationality and rigorousness.
@@
Appendix G Related Work
G.1 LLM-as-a-Judge
@@
In recent years, Large Language Models (LLMs) have emerged as a transformative technology, garnering global attention and stimulating substantial research into their applications. For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature [1]. However, it lacks scalability and reproducibility [2].
@@
As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks [3, 4, 21, 10, 22, 23, 24, 25, 26, 11, 1, 27, 28, 29, 30, 2, 31, 17, 32, 33, 34, 35, 36, 37, 38, 7, 12, 39], including open-ended story generation [30], adversarial attacks [21], summarization [2], machine translation [29], and instruction following [1].
@@
These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles [3, 40, 41, 21, 28, 10, 6, 22, 42, 5, 43, 44, 45, 11, 1, 8, 17, 39, 7, 46, 12, 18, 9, 19]. As LLMs have made content generation significantly easier, the volume of generated responses has increased, making it impractical to rely solely on human evaluation. Therefore, cost-effective LLM judges are needed to assess these responses efficiently. LLM-as-a-Judge is typically employed in Q&A evaluation tasks, where the LLM judge is prompted to evaluate the quality of responses, usually generated by models, answering the questions.
@@
In many of these tasks, LLM judges have shown a high level of agreement with human evaluators [3, 4, 5], yet in some tasks, they are less effective, largely due to inherent biases [28, 21, 40]. Even in cases where agreement is high, judgments may still suffer from biases.
@@
When employing LLM-as-a-Judge, various types of judging are available, which can be categorized either by the scale or comparative method. From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment [6]. By comparative method, it can be either score-based or relation-based [7].
@@
For example, pointwise scoring [17] lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics. Pairwise/Listwise Scoring [2, 5, 19, 3] prompts the LLM judges to score a pair/list of model-generated answers. Listwise Ranking is another relation-based assessment that, instead of giving a score, requires the LLM judge to rank a list of responses following some specified order (e.g., from best to worst). Pairwise comparative assessment [2, 19, 3], on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
@@
G.2 Position Bias
Previous studies have discovered and investigated multiple types of biases, such as position bias [3, 4, 19, 22, 35, 21, 8, 5, 34, 40, 36, 1, 47, 6, 45, 46, 10, 11, 17, 2, 20, 25, 41, 28, 30, 29, 44, 12, 7, 43], verbosity/length bias [3, 11, 17], self-enhancement bias [12, 3], selection bias [40, 41], and contextual bias [11, 48].
Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source [39, 18], proprietary commercial ones [13, 14, 15], and fine-tuned models [42, 47, 5, 39, 17].
To clarify, the aforementioned and following position bias refers to the concept within the context of LLM-as-a-Judge, meaning that LLM judge tends to favor responses based on their position in the prompt rather than their content. For example, in a pairwise comparative assessment scenario, if the LLM judge consistently selects the first response as superior even after switching the order of the two responses (same position, but different content), then a position bias occurs. Some other studies also use the term position bias [49, 50, 51, 52], but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
Position bias is arguably the most prevalent and impactful bias among all. [9] notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought [53] performance over self-training controls for all biases except position bias. Furthermore, Khan et al. [8] point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
Moreover, there is ongoing debate over whether selection bias originates from position bias. [41] argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity.[43] contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
@@ G.2 Position Bias
G.3 Deal with Position Bias
Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts [3, 43, 41, 54, 55]. Also, LLM judges are vulnerable to attacks [43, 44, 46].
Parse data with position bias There are many ways to deal with the position bias. The naive way is to exclude the inconsistent judgments(i.e. if the LLM judge gives a positionally biased judgment on a pair or list of model responses) [3, 44, 45, 39]. While this ensures consistent and reliable remaining judgments, it does not resolve the fundamental problem. Moreover, if the LLM judge is highly biased, this method discards valuable evaluation instances and information, making it an ineffective and somewhat desperate measure.
To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging [3, 43, 39, 45] or regard inconsistency as a half-win or tie for relation-based judging [3, 12] after swapping the order of model responses in the prompt. For instance, in a pairwise scoring scenario, if model A receives a score of 8 when put in the first position and 4 when put in the second position, the overall score for it compared to model B would be 6; in the pairwise comparative case, if model A wins when put in the first position but then lose when put in the second position, it counts as a tie or half-win for both model A and model B. The latter way of swapping + tie is proposed because intuitively the position bias is more likely to occur when the model responses get evaluated share a similar quality in terms of the evaluating metric. Our quantitative study also verifies this intuition, evidenced by the fact the LLM judge’s positional consistency (the percentage of positionally consistent pairs of judgment) is positively proportional to the answer quality gap between the model responses.
@@ G.3 Deal with Position Bias
Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping [20], split-and-merge [7], and multi-agent discussion [12, 8]. However, these methods are either costly and time-consuming (e.g., multi-agent discussion and review to reach agreement) or ineffective.
Furthermore,[11] suggests that existing calibration techniques designed to reduce bias, including context decomposition [7], order permutation [45, 40], ensembling [56], and batch calibration [48], are insufficient to align LLM evaluators, even with supervised data.
@@
For instance, [7] proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%. However, the extraordinarily high consistency of the original evaluation may imply that the choice of answer pairs may be biased in terms of answer quality gap, a factor that our study proves to be significantly impactful. In other words, if the quality between the two model responses differs considerably, the LLM judges will make little position bias during judgment and hence is easy to calibrate. This example illustrates how lacking a comprehensive understanding of the factors affecting position bias can lead to overestimating or misjudging the effectiveness of methods proposed to resolve it and improve the performance of LLM-as-a-Judge.
@@
While prior research includes some positional preference results [3, 45, 5, 12, 10], none have conducted a detailed analysis. They focus primarily on positional consistency in relation to position bias, while we regard positional preference as an equally crucial component that needs to be examined, understood, and enhanced in future work.
@@
Apart from that, previous studies have yet to examine how stable LLM judges are across repetitions. MLLM-as-a-Judge [19] conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments. In comparison, we investigate the repetition stability to determine whether position bias is solely influenced by the prompt’s positional information and the judge’s intrinsic properties or can be partly attributed to judgment randomness. If the judgments exhibit randomness over repetitions, then the position bias overlaps with repetition bias, complicating the issue and potentially invalidating previous findings.
G.4 Summary of Prior Work